The compensated n-type silicon at 300 K has a conductivity: 16 (ohm-cm)-1 and an acceptor doping concentration: 1017 cm-3.

a) Express the mobility as a function of the doping concentration.

b) Calculate the mobility and conductivity when Nd=2 1017 cm-3.
Answer #1

is Nd= Mobility as in function of doping concentration or donor concentration accepto en doping concentrecction. Let assume cMobility ) 6 = e ( Nd - Na) 6 = 16 (s2.cm- 2 x 1017 cm 3 Nd = Na = 1017 cm 3 ein 16 (1.6x10-19) [(@x1017) – 1017] at en = 100